OpenStudy (ksaimouli):

an object dropped from rest from the top of a tall building falls y=16t^2 feet first t seconds (a)find teh speed of a object at t=3s and confirm algebraically

OpenStudy (ksaimouli):

@Chyme u r right could you continue from the lim plz

OpenStudy (anonymous):

lim Δx→0 f(x+Δx)−f(x) /Δx in this case we substitute f(x) for f(t) and we get \[\lim Δt→0\frac{ f(t+Δt)−f(t)}{ Δt } \] Now plug in f(t)=16t^2 \[\lim Δt→0\frac{ 16(t+\Delta t)^2−(16t^2)}{ Δt } \]

OpenStudy (anonymous):

with me so far?

OpenStudy (ksaimouli):

yes

OpenStudy (anonymous):

wanna take a shot at simplifying and solving?

OpenStudy (ksaimouli):

yes

OpenStudy (anonymous):

you want to make sure you get rid of the Delta t on the bottom by some algebra

OpenStudy (ksaimouli):

alright thx

OpenStudy (anonymous):

np